# Application of the Lovász-Schrijver Lift-and-Project Operator to Compact Stable Set Integer Programs

### v2

@@ Abstract
The Lovász theta function $\theta(G)$ provides a very good upper bound on the stability number of a graph $G$. It can be computed in polynomial time by solving a semidefinite program (SDP), which also turns out to be fairly tractable in practice. Consequently, $\theta(G)$ achieves a hard-to-beat trade-off between computational effort and strength of the bound. Indeed, several attempts to improve the theta bound are documented, mainly based on playing around the application of the $N_{+}(\cdot)$ lifting operator of Lovász and Schrijver to the classical formulation of the maximum stable set problem. Experience shows that solving such SDP-s often struggles against practical intractability and requires highly specialized methods. We investigate the application of such an operator to two different linear formulations based on clique and nodal inequalities, respectively. Fewer inequalities describe these two and yet guarantee that the resulting SDP bound is at least as strong as $\theta(G)$. Our computational experience, including larger graphs than those previously documented, shows that upper bounds stronger than $\theta(G)$ can be accessed by a reasonable additional effort using the clique-based formulation on sparse graphs and the nodal-based one on dense graphs.
@@ 1 Introduction
Given a simple undirected graph $G=(V,E)$, with vertex set $V$ and edge set $E$, a stable (or independent) set in $G$ is a set of pairwise non-adjacent vertices. A fundamental problem in combinatorial optimization is the Stable Set Problem (SSP), which consists of computing a stable set of maximum cardinality [33, 50]. This value, denoted by $\alpha(G)$, is the stability number of $G$. A weighted version of this problem can be considered if a weight vector $w\in\mathbb{Q}_{+}^{n}$ is given. SSP is NP-hard in the strong sense and even hard to approximate [35], it is equivalent to the Maximum Clique and Set Packing problems [53] and has a wide range of applications in practical contexts [10, 11, 16, 14, 57, 68]. SSP is naturally formulated as a binary program by the so-called edge inequalities:
$\alpha_{w}(G)=\max$ $\sum_{i\in V}w_{i}x_{i}$
@@
denotes the stable set polytope, that is, the convex hull of the incidence vectors of all stable sets in $G$. The intense study of STAB($G$) has led, since the early seventies, to the identification of several classes of valid inequalities, often associated with specific “facet-producing” sub-graphs [33, 11, 2, 15]. In practice, however, linear relaxations based on these inequalities often yielded rather weak upper bounds on the stability number of unstructured graphs. As a consequence, basic branch-and-cut algorithms are not particularly effective [49] and much more sophisticated cut generation procedures [30, 31, 20, 19, 56, 55] are required to reduce the performance gap with fast combinatorial exact algorithms, such as those devised in [58, 52, 62].
@@
In his seminal paper, Lovász [42] introduced the celebrated theta number of a graph, denoted by $\theta(G)$. It represents an upper bound for $\alpha(G)$, which can be computed in polynomial time (up to arbitrary precision) by solving a Semidefinite Program (SDP) [33]. Since its introduction, $\theta(G)$, along with the associated non-polyhedral convex theta body $\operatorname{TH}(G)$ [33], turned out to be powerful tools. For instance, they allowed to prove that the SSP can be solved in polynomial time when $G$ is a perfect graph [33]. Indeed, the practical relevance of $\theta(G)$ is no less. On one side, it often turns out to be a significantly stronger upper bound than those from linear relaxations [28, 39]. On the other side, the structure of the associated SDP relaxation often mitigates numerical difficulties that typically affect SDP algorithms. As a consequence, $\theta(G)$ realizes a very good trade-off between the quality of the upper bound and computational effort, which makes it attractive to be incorporated within branch-and-bound algorithms, as studied by, e.g., Wilson [67] and more recently by Gaar et al. [25].
@@
In a successive landmark paper Lovász and Schrijver [43] introduced the $N_{+}(\cdot)$ lift-and-project operator, which can be applied to the LP relaxation of any 0-1 LP and returns a stronger SDP relaxation. Its application to $\operatorname{FRAC}(G)$ has also been intensely investigated and has resulted tighter than the Lovász theta relaxation. From a theoretical perspective, its strength has been documented by showing implications of (additional) facet defining inequalities of the stable set polytope [43, 27]. Moreover, as noted in [7, 8], it provides an effective tool to identify other classes of graphs for which the SSP can be solved in polynomial time. Recently, graph classes yielding stable set polytopes with high rank with respect to $N_{+}(\cdot)$ have been studied in [1].
@@
From the computational point of view, things are somewhat more involved. Experiments conducted with $N_{+}(\operatorname{FRAC}(G))$ [21, 12] and related variants [23, 34] show that these relaxations may yield stronger upper bounds than $\theta(G)$ but at the price of considerably increasing the computational effort. In fact, solving these large-scale SDPs requires specialized methods and large instances can hardly be managed. Other methods, based on strengthening SDP relaxations by linear inequalities not related to the $N_{+}(\cdot)$ operator, have also been presented by Gaar and Rendl [24] and Locatelli [41]. These achieve strong bounds but still require a significant computational effort.
@@ 2 Preliminaries
This paper will adopt the following standard notation. The set of integers $\{1,...,k\}$ is denoted by $[k]$. The complement of a graph $G=(V,E)$ is the graph ${\bar{G}}=\big{(}V,{\bar{E}}\big{)}$, where ${\bar{E}}=\big{\{}\{i,j\}\subset V:\,\{i,j\}\in(V\times V)\setminus E\big{\}}$. Given a vertex $i\in V$, we let $\Gamma_{G}(i)$ denote the set of neighbours of $i$ in $G$, i.e., $\Gamma_{G}(i)=\big{\{}j\in V\setminus\{i\}:\{i,j\}\in E\big{\}}$. When no confusion arises, we let $\Gamma_{G}(i)=\Gamma(i)$. Finally, given a vertex $S\subseteq V$, we let $G[S]=\big{(}S,E(S)\big{)}$ denote the subgraph of $G$ induced by $S$, and $r(S)$ denote its stability number $\alpha\big{(}G[S]\big{)}$ (the so-called rank of $S$). We also let ${\cal S}_{n}$ be the set of real symmetric square matrices of order $n$ and ${\cal S}^{+}_{n}~{}\subset~{}{\cal S}_{n}$ denotes the cone of those which are positive semidefinite (PSD). PSDness of a matrix $Y\in{\cal S}^{+}_{n}$ is also denoted by $Y\succeq 0$. Given any square matrix $X\in\mathbb{R}^{n\times n}$ we denote by $(X)_{ij}$ the element at the $i$-th row and the $j$-th column, by $\operatorname{diag}(X)\in\mathbb{R}^{n}$ its main diagonal and by $\operatorname{rank}(X)$ its rank. We denote by $\left\langle X,Y\right\rangle=\operatorname{tr}(XY)$ the standard inner product in ${\cal S}_{n}$, where $\operatorname{tr}(A)$ is the trace of the square matrix $A$. At last, we denote with $J$ the all-ones square matrix and $e$ the all-ones vector; the dimensions should be clear from the context where not explicit.
@@ 2.1 The lift-and-project operator of Lovász and Schrijver
$\text{max}\left\{c^{\top}x\ |\ Ax\leq b,\ x_{i}\in\{0,1\}\text{ for }i\in I=\{1,\ldots,n\}\right\},$ (IP)
@@
$x_{i}((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m],$ (2)
$(1-x_{i})((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m].$ (3)
which are valid for $P$ as $x_{i}\geq 0$ and $1-x_{i}\geq 0$ for all $x\in P$. These inequalities can be linearized through the introduction of the matrix variable $X=xx^{\top}$ and rewriting (2) and (3) as:
@@
${N_{+}(P)}=\left\{x\in\mathbb{R}^{n}\ \Big{|}\ \exists\ Y\in M_{+}(\text{P})\text{ with }\begin{pmatrix}1\\ x\end{pmatrix}=\text{diag}(Y)\right\}.$
[…]
@@
Moreover, Bodur et al. [9] noted that the definition of $M_{+}(P)$ and $N_{+}(P)$ does not depend on the representation of $P$ but only on the set of points in $P$. Hence, the following property holds:
@@
Given two polytopes $P,P^{\prime}$ with $P^{\prime}\subseteq P$ then ${M_{+}(P^{\prime})}\subseteq{M_{+}(P)}$
@@ 3.1 Linear relaxations of the SSP
Several classes of valid inequalities for $\operatorname{STAB}(G)$ have been identified since early seventies [33, 11]. Among them, clique inequalities $\sum_{i\in C}x_{i}\leq 1$, where $C\subseteq V$ is a maximal clique (set of pairwise adjacent vertices), induce facets of STAB$(G)$ [53]. As customary in the literature, we define the polytope
@@
where ${\cal K}$ denotes the collection of all maximal cliques of $G$. As clique inequalities clearly imply edge inequalities, STAB$(G)\subseteq$ $\operatorname{QSTAB}(G)$subseteq FRAC$(G)$ and inclusion is generally strict [33]. Clique inequalities in $G$ can be exponential in number, and the associated separation problem is NP-hard in the strong sense. However, effective separation heuristics make this class of inequalities computationally easy to handle, as shown in [45].
The 0-1 linear program
$\max$ $\sum\limits_{i\in V}w_{i}x_{i}$
s.t. $\sum\limits_{i\in C}x_{i}\leq\,1\quad(C\in{\mathcal{C}})$ (8)
$x_{i}\in\{0,1\}\quad(i\in V).$
is a valid formulation for the SSP if ${\mathcal{C}}\subseteq{\cal K}$ is any collection of maximal cliques covering all the edges of $G$ (i.e., both the endpoints of every edge are contained in at least one clique of $\mathcal{C}$). Extending the standard notation, we denote by $\operatorname{QSTAB}(G,\mathcal{C})$ the polytope associated with its continuous relaxation. It’s not hard to see that ${\operatorname{QSTAB}(G)}\subseteq{\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC(G)}$. $\operatorname{QSTAB}(G,\mathcal{C})$ has a number of clique inequalities bounded by $|E|$ and it typically contains considerably fewer inequalities than $\operatorname{FRAC(G)}$. In fact, formulations based on a greedily computed minimum size collection $\mathcal{C}$ are competitive in practice, as shown in several studies (see, e.g., Letchford et al. [40] and references therein).
Other rank inequalities of the form $\sum_{i\in S}x_{i}\leq\alpha(G[S])$ have been studied for vertex subsets $S\subseteq V$ inducing special graphs (see e.g., Borndörfer [11] and Giandomenico et al. [28] for a survey). A well-known example arises when $S=H$, $|H|\geq 5$, induces a chordless cycle of odd cardinality, yielding the odd-hole inequality $\sum_{i\in H}x_{i}\leq\lfloor\frac{|H|}{2}\rfloor$. Among other subgraphs, we recall webs and antiwebs introduced in [63]. Let $p$ and $q$ be integers satisfying $p>2q+1$ and $q>1$, and use arithmetic modulo $p$. A $(p,q)$--web is a graph with vertex set $\{1,...,p\}$ and with edges from $i$ to $\{i+q,\ldots,i-q\}$, for every $1\leq i\leq p$. A $(p,q)$--antiweb is the complement of a $(p,q)$--web. The web inequalities take the form $\sum_{i\in W}x_{i}\leq q$ for every vertex set $W$ inducing a $(p,q)$-web, and the antiweb inequalities take the form $\sum_{i\in AW}x_{i}\lfloor p/q\rfloor$ for every vertex set $AW$ inducing a $(p,q)$-antiweb.
A different modeling rationale has been pursued by Della Croce and Tadei [22] who considered the surrogate relaxation of formulation (1) obtained by summing up, for every $i\in V$, the edge inequalities over all $j\in\Gamma(i)$. This yields the so called nodal inequality $\sum_{j\in\Gamma(i)}x_{j}+|\Gamma(i)|x_{i}\,\leq\,|\Gamma(i)|$. Murray and Church [48] observed that this relaxation can be strengthened by replacing $|\Gamma(i)|$ with any value $r_{i}\geq\alpha(G[\Gamma(i)])$, which returns a class of 0-1 linear programs of the form
$\max$ $\,\,\sum\limits_{i\in V}w_{i}x_{i}$
s.t. $\sum\limits_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i}\quad(i\in V)$ (9)
$x_{i}\in\{0,1\}\quad(i\in V).$
@@
${\operatorname{NOD}(G,r)}=\left\{x\in[0,1]^{|V|}:\sum_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i},\hskip 5.0pti\in V\right\}$
to denote the polytope defined from (9) (and non-negativity), where $r\in\{\Gamma,\theta,\alpha\}$ specifies the coefficient used. Clearly, we have ${\operatorname{NOD}(G,\alpha)}\subseteq{\operatorname{NOD}(G,\theta)}\subseteq{\operatorname{NOD}(G,\Gamma)}$. It is worth noting that nodal inequalities in general do not imply edge inequalities, even in the case $r=\alpha$. On the other hand, if $r=\alpha$, they imply the wheel inequalities (see Borndörfer [11] for their definition) which are not implied by clique inequalities.

### v3

# Application of the Lovász-Schrijver Operator to Compact Stable Set Integer Programs

@@ Abstract
The Lovász theta function $\theta(G)$ provides a very good upper bound on the stability number of a graph $G$. It can be computed in polynomial time by solving a semidefinite program (SDP), which also turns out to be fairly tractable in practice. Consequently, $\theta(G)$ achieves a hard-to-beat trade-off between computational effort and strength of the bound. Indeed, several attempts to improve the theta bound are documented, mainly based on playing around the application of the $N_{+}(\cdot)$ lifting operator of Lovász and Schrijver to the classical formulation of the maximum stable set problem (SSP). Experience shows that solving such SDPs often struggles against practical intractability and requires highly specialized methods. We investigate the application of such an operator to two different linear formulations of the SSP based on clique and nodal inequalities, respectively. These two formulations are described by fewer inequalities than the natural formulation based on edge inequalities, yet they guarantee that the resulting SDP bound is at least as strong as $\theta(G)$. Our computational experience, including larger graphs than those previously documented, shows that upper bounds stronger than $\theta(G)$ can be accessed by a reasonable additional effort using the clique-based formulation on sparse graphs and the nodal-based one on dense graphs.
@@ 1 Introduction
Given a simple undirected graph $G=(V,E)$, with vertex set $V$ and edge set $E$, a stable (or independent) set in $G$ is a set of pairwise non-adjacent vertices. A fundamental problem in combinatorial optimization is the Stable Set Problem (SSP), which consists of computing a stable set of maximum cardinality [33, 50]. This value, denoted by $\alpha(G)$, is the stability number of $G$. For the sake of presentation, we do not explicitly address the weighted version of this problem, where a weight vector $w\in\mathbb{Q}_{+}^{|V|}$ is given, although all the results presented remain valid in that setting. The SSP is NP-hard in the strong sense and even hard to approximate [35], it is equivalent to the Maximum Clique and Set Packing problems [53] and has a wide range of applications in different practical contexts, see, e.g., those illustrated in [14, 16, 57] and those surveyed in [10, 68]. The SSP is naturally formulated as a binary program by the so-called edge inequalities
$\alpha(G)=\max$ $\sum_{i\in V}x_{i}$
@@
denotes the stable set polytope, that is, the convex hull of the incidence vectors of all stable sets in $G$. The intense study of STAB($G$) has led, since the early seventies [2], to the identification of several classes of valid inequalities, often associated with specific “facet-producing” sub-graphs [11, 15, 33]. In practice, however, linear relaxations based on these inequalities often yielded rather weak upper bounds on the stability number of unstructured graphs. As a consequence, basic branch-and-cut algorithms are not particularly effective [49] and much more sophisticated cut generation procedures, see [19, 20, 30, 31, 55, 56], are required to reduce the performance gap with fast combinatorial exact algorithms, such as those devised in [52, 58, 62].
@@
In his seminal paper, Lovász [42] introduced the celebrated theta number of a graph, denoted by $\theta(G)$. It represents an upper bound on $\alpha(G)$, which can be computed in polynomial time (up to arbitrary precision) by solving a Semidefinite Program (SDP). Since its introduction, $\theta(G)$, along with the associated non-polyhedral convex theta body $\operatorname{TH}(G)$, turned out to be powerful tools. For instance, they allowed to prove that the SSP can be solved in polynomial time when $G$ is a perfect graph. For the description of these results we refer the reader to the book of Grötschel et al. [33]. Indeed, the practical relevance of $\theta(G)$ is no less. On one side, it often turns out to be a significantly stronger upper bound than those from linear relaxations, as documented in [28, 39]. On the other side, the structure of the associated SDP relaxation often mitigates numerical difficulties that typically affects SDP algorithms. As a consequence, $\theta(G)$ realizes a very good trade-off between the quality of the upper bound and computational effort, which makes it attractive to incorporate it within branch-and-bound algorithms, as studied by, e.g., Wilson [67] and more recently by Gaar et al. [25].
@@
In a successive landmark paper Lovász and Schrijver [43] introduced the $N_{+}(\cdot)$ lift-and-project operator, which can be applied to the LP relaxation of any 0-1 LP and returns a stronger SDP relaxation. Its application to $\operatorname{FRAC}(G)$ has also been intensively investigated and has resulted in a tighter relaxation than the Lovász theta relaxation. From a theoretical perspective, its strength has been documented by showing implications of (additional) facet defining inequalities of the stable set polytope (Giandomenico and Letchford [27], Lovász and Schrijver [43]). Moreover, Bianchi et al. [7, 8] showed that it provides an effective tool to identify other classes of graphs for which the SSP can be solved in polynomial time. Recently, graph classes yielding stable set polytopes with high rank with respect to $N_{+}(\cdot)$ have been studied in [1].
@@
From the computational point of view, things are somewhat more involved. Experiments conducted with $N_{+}(\operatorname{FRAC}(G))$ by Dash [21], Burer and Vandenbussche [12] and with related variants by Dukanovic and Rendl [23], Gruber and Rendl [34], show that these relaxations may yield stronger upper bounds than $\theta(G)$ but at the price of considerably increasing the computational effort. In fact, solving these large-scale SDPs requires specialized methods, and large instances can hardly be managed. Other methods, based on strengthening SDP relaxations by linear inequalities not related to the $N_{+}(\cdot)$ operator, have also been presented by Gaar and Rendl [24] and Locatelli [41]. These achieve strong bounds but still require significant computational effort.
@@ 2 Preliminaries
This paper will adopt the following standard notation. The set of integers $\{1,...,k\}$ is denoted by $[k]$. The complement of a simple graph $G=(V,E)$ is the graph ${\bar{G}}=\big(V,{\bar{E}}\big)$, where ${\bar{E}}=\big\{\{i,j\}\subset V:\,\{i,j\}\in(V\times V)\setminus E,i\not=j\big\}$. Given a vertex $i\in V$, we let $\Gamma_{G}(i)$ denote the set of neighbors of $i$ in $G$, i.e., $\Gamma_{G}(i)=\big\{j\in V\setminus\{i\}:\{i,j\}\in E\big\}$. When no confusion arises, we let $\Gamma_{G}(i)=\Gamma(i)$. Finally, given a set $S\subseteq V$, we let $G[S]=\big(S,E(S)\big)$ denote the subgraph of $G$ induced by $S$, and $r(S)$ denote its stability number $\alpha\big(G[S]\big)$ (the so-called rank of $S$). We also let ${\cal S}_{n}$ be the set of real symmetric square matrices of order $n$ and ${\cal S}^{+}_{n}~\subset~{\cal S}_{n}$ denotes the cone of those which are positive semidefinite (PSD). PSDness of a matrix $Y\in{\cal S}^{+}_{n}$ is also denoted by $Y\succeq 0$. Given any square matrix $X\in\mathbb{R}^{n\times n}$ we denote by $X_{ij}$ the element at the $i$-th row and the $j$-th column, by $\operatorname{diag}(X)\in\mathbb{R}^{n}$ its main diagonal and by $\operatorname{rank}(X)$ its rank. We denote by $\left\langle X,Y\right\rangle=\operatorname{tr}(XY)$ the standard inner product in ${\cal S}_{n}$, where $\operatorname{tr}(A)$ is the trace of the square matrix $A$. At last, we denote by $J$ the all-ones square matrix and $e$ the all-ones vector; the dimensions should be clear from the context where not explicit. Given $a\in\mathbb{R}^{n}$ and $b\in\mathbb{R}$, the linear inequality $a^{\top}x\leq b$ is valid for $\operatorname{STAB}(G)$ if $\left\{x\in\mathbb{R}^{n}\mid a^{\top}x\leq b\right\}\supseteq\operatorname{STAB}(G).$
@@ 2.1 The lift-and-project operator of Lovász and Schrijver
$\text{max}\left\{c^{\top}x\ |\ Ax\leq b,\ x_{i}\in\{0,1\}\text{ for }i\in[n]\right\},$ (IP)
@@
$x_{i}\cdot((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m],$ (2)
$(1-x_{i})\cdot((a^{k})^{\top}x-b^{k})$ $\leq 0\quad i\in[n],k\in[m].$ (3)
which are valid for $P$ as $x_{i}\geq 0$ and $1-x_{i}\geq 0$ for all $x\in P$. These inequalities can be linearized through the introduction of the matrix variable $X=xx^{\top}$ and  (2) and (3) can be rewritten as
@@
${N_{+}(P)}=\left\{x\in\mathbb{R}^{n}\ \Big|\ \exists\ Y\in M_{+}(P)\text{ with }\begin{pmatrix}1\\ x\end{pmatrix}=\text{diag}(Y)\right\}.$
Lovász and Schrijver [43] showed that $N_{+}(P)$ is contained in $P$ and is a valid relaxation for $P_{I}$, i.e. $P_{I}\subseteq{N_{+}(P)}\subseteq P$ holds. For the ease of presentation, in the reminder of this paper we will refer to both $M_{+}(P)$ and $N_{+}(P)$ as a relaxation of $P_{I}$, despite the former “lives” in the lifted variables space. Furthermore, note that all the linear inequalities $Ax\leq b$ are trivially satisfied by points in $M_{+}(P)$.
@@
Moreover, Bodur et al. [9] noted that the definition of $M_{+}(P)$ and $N_{+}(P)$ does not depend on the representation of $P$ but only on the set of points in $P$. Hence, the following property holds
@@
Given two polytopes $P,P^{\prime}$ with $P^{\prime}\subseteq P$ then ${M_{+}(P^{\prime})}\subseteq{M_{+}(P)}$.
@@ 3.1 Linear relaxations of the SSP
Several classes of valid inequalities for $\operatorname{STAB}(G)$ have been identified since early seventies, as is illustrated in [11, 33]. Among them, clique inequalities $\sum_{i\in C}x_{i}\leq 1$, where $C\subseteq V$ is a maximal clique (set of pairwise adjacent vertices), induce facets of STAB(G)(G) [53]. As customary in the literature, we define the polytope
@@
[…]
$\max$ $\sum\limits_{i\in V}x_{i}$
s.t. $\sum\limits_{i\in C}x_{i}\leq\,1\quad C\in{\mathcal{C}}$ (8)
$x_{i}\in\{0,1\}\quad i\in V.$
is a valid formulation for the SSP , for all ${\mathcal{C}}\in{\cal Q}$. Extending the standard notation, we denote by $\operatorname{QSTAB}(G,\mathcal{C})$ the polytope associated with its continuous relaxation. Unless stated otherwise, the results presented in the remainder of the paper are valid for any ${\mathcal{C}}\in{\cal Q}$. It is not hard to see that ${\operatorname{QSTAB}(G)}\subseteq{\operatorname{QSTAB}(G,\mathcal{C})}\subseteq\operatorname{FRAC}(G)$. If ${\mathcal{C}}\in{\cal Q}$ is a minimal edge clique cover, then $|\mathcal{C}|\leq|E|$.
In practice, formulations based on greedily computed edge clique covers $\mathcal{C}$, which are made up of maximal cliques, contain considerably fewer inequalities than those describing $\operatorname{FRAC(G)}$, and also provide a good approximation of $\operatorname{QSTAB}(G)$}, as shown in several studies (see, e.g., Letchford et al. [40] and references therein).
Other rank inequalities of the form $\sum_{i\in S}x_{i}\leq\alpha(G[S])$ have been studied for vertex subsets $S\subseteq V$ inducing special graphs (see e.g., Borndörfer [11] and Giandomenico et al. [28] for a survey). A well-known example arises when $S=H$, $|H|\geq 5$, induces a chordless cycle of odd cardinality, yielding the odd-hole inequality $\sum_{i\in H}x_{i}\leq\lfloor|H|/2\rfloor$. Among other subgraphs, we recall webs and antiwebs introduced by Trotter Jr [63]. Let $p$ and $q$ be integers satisfying $p>2q+1$ and $q>1$, and use arithmetic modulo $p$. A $(p,q)$-web is a graph with vertex set $\{1,...,p\}$ and with edges from $i$ to $\{i+q,\ldots,i-q\}$, for every $1\leq i\leq p$. A $(p,q)$-antiweb is the complement of a $(p,q)$-web. The web inequalities take the form $\sum_{i\in W}x_{i}\leq q$ for every vertex set $W$ inducing a $(p,q)$-web, and the antiweb inequalities take the form $\sum_{i\in AW}x_{i}\leq\lfloor p/q\rfloor$ for every vertex set $AW$ inducing a $(p,q)$-antiweb.
A different modeling rationale has been pursued by Della Croce and Tadei [22] who considered the surrogate relaxation of formulation based on (1) obtained by summing up, for every $i\in V$, the edge inequalities over all $j\in\Gamma(i)$. This yields the so called nodal inequality $\sum_{j\in\Gamma(i)}x_{j}+|\Gamma(i)|x_{i}\,\leq\,|\Gamma(i)|$. Murray and Church [48] observed that this relaxation can be strengthened by replacing $|\Gamma(i)|$ with any value $r_{i}\geq\alpha(G[\Gamma(i)])$, which returns a class of 0-1 linear programs of the form
$\max$ $\,\,\sum\limits_{i\in V}x_{i}$
s.t. $\sum\limits_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i}\quad i\in V$ (9)
$x_{i}\in\{0,1\}\quad i\in V.$
@@
${\operatorname{NOD}(G,r)}=\left\{x\in[0,1]^{|V|}:\sum_{j\in\Gamma(i)}x_{j}+\,r_{i}\,x_{i}\,\leq\,r_{i},\enskip i\in V\right\}$
to denote the polytope defined from (9) (and non-negativity), where $r\in\{\Gamma,\theta,\alpha\}$ specifies the coefficient used. Clearly, we have
${\operatorname{NOD}(G,\alpha)}\subseteq{\operatorname{NOD}(G,\theta)}\subseteq{\operatorname{NOD}(G,\Gamma)}.$
It is worth noting that nodal inequalities in general do not imply edge inequalities, even in the case $r=\alpha$. On the other hand, if $r=\alpha$, they imply the wheel inequalities (see Borndörfer [11] for their definition) which are not implied by clique inequalities.